OpenStudy (anonymous):

solve the equation x-1/3=3

OpenStudy (angela210793):

x-1/3=3(+1/3 both sides) x=3+1/3 x=10/3

OpenStudy (anonymous):

do you mean (x-1)/3 or x -(1/3)?

OpenStudy (anonymous):

x - (1/3) = 3 x = 3 + (1/3) x = 10/3

OpenStudy (anonymous):

@virtue its the way typed it is the way i meant it.. thanks